OpenStudy (anonymous):

Write the expression as a single log: log x-log(x+3)+log(x^2-9)

OpenStudy (anonymous):

log x-log(x+3)+log(x^2-9) loa+logb+logc-logd-loge \[=\log \frac{ a~b~c }{ d~e }\]

OpenStudy (raden):

ca we cancels the similar terms ? :v